OpenStudy (anonymous):

9x^3+25x=30x^2 Solve Solving Equations by Factoring

OpenStudy (anonymous):

set all terms on one side or set the equation to 0. 9x^3+25x-30x^2 = 0 factor out an x. x(9x^2-30x+25) = 0 one of the x values = 0 (x)=0 (9x^2-30x+25)=0 factor by unfoiling (9x^2-30x+25)=0 (3x-5)(3x-5) = 0 add 5 to both sides 3x=5 solving for x will get you: x = (5/3) so the zeros of this equation are (this is the answer if they're asking you to find the zeros): x = 0 and x = (5/3)